# Magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances in hot and dense strange hadronic matter

### v1

@@ IV.2 In-medium magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances
Figure 6: Effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
@@
Figure 7: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
@@
Figure 8: Effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
@@
Figure 9: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
@@
Figure 10: Effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
@@
Figure 11: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
@@
Figure 12: Effective magnetic moments of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
@@
Figure 13: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
@@
In this section we elaborately discuss the results of the vacuum and in-medium magnetic moments of ${\frac{1}{2}}^{-}$ octet baryon resonances $p^{*},n^{*},\Sigma^{*+},\Sigma^{*-},\Sigma^{*0},\Xi^{*-}$, $\Xi^{*0}$ and low-lying $N^{*}$ resonances ${S^{+}_{11}}(1650)$ and ${S^{0}_{11}}(1650)$, observed from figs. 6, 7, 8, 9, 10, 11, 12 and 13. For the sake of discussion, we have also included the individual contributions from valence and sea quarks to the total baryonic magnetic moment which allows us to better interpret the results.
@@
In fig. 6, we have shown the magnetic moment of $p^{*}$ and $n^{*}$ resonances with density of the medium at temperatures $T=0$ MeV and $T=100$ MeV. The observed in-medium behavior of $\mu^{*}_{p^{*}}$ and $\mu^{*}_{n^{*}}$, shows a steeper slope at lower densities of nuclear matter ($\rho_{B}\leq 2\rho_{0}$) and gradually attains saturation at much higher densities. As we trace the plot of these effective baryon magnetic moments while the medium density rises to $\rho_{B}=3\rho_{0}$, we observe that $\mu^{*}_{p^{*}}$ increases by $0.59\mu_{N}$ and $\mu^{*}_{n^{*}}$ initially approaches a minima at $-1.382\mu_{N}$ (at $\rho_{B}\approx\rho_{0}$) and then climbs up close to its vacuum value at $-1.375\mu_{N}$. A conclusive understanding of the observed effective baryonic magnetic moments of $p^{*}$ and $n^{*}$ can be provided by analyzing fig. 7, where the contributions from valence quarks, sea quarks and orbital angular momentum of the quark sea have been shown individually. For example, in case of $n^{*}$ in symmetric nuclear matter, the decrease in the valence quark contributions dominates over the increasing contributions of sea quarks and the orbital contributions of sea quarks, which leads to an overall decreasing behavior of $\mu^{*}_{n^{*}}$ below the saturation density of the medium ($\rho_{0}$). As the medium density moves past this point ($\rho_{B}\approx\rho_{0}$), owing to the dominance of $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$ contributions the effective in-medium behavior of $\mu^{*}_{n^{*}}$ reverses and it approaches higher values. On a similar note, the density dependency of $\mu^{*}_{p^{*}}$ can be comprehended by taking into account the increase in $\mu^{*}_{val}$ dominating over the decrease in $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$. In symmetric strange matter at $f_{s}=0.3$, the in-medium behavior of $\mu^{*}_{p^{*}}$ and $\mu^{*}_{n^{*}}$ is enhanced. From fig. 6 we can infer that, the finite strangeness fraction in the medium leads to a splitting in baryonic magnetic moments of $p^{*}$ and $n^{*}$, which is visible beyond the nuclear saturation density, $\rho_{0}$. All other medium parameters affixed, non-zero isospin asymmetry ($i_{a}=0,0.3,0.5$) causes fine splitting in the $\mu^{*}_{p^{*}}$ and $\mu^{*}_{n^{*}}$ curves as can be seen from fig. 6, which has also been identified in their respective $\mu^{*}_{val}$, $\mu^{*}_{sea}$, $\mu^{*}_{orbit}$ contributions by referring to fig. 7. As we have included “Goldstone Boson Exchange”(GBE) effect in this study, it is important to note that the sea quark effect has an opposite sign as compared to the valence quark effect, whereas the orbital angular momentum of sea quarks has the same contribution as that from the valence quark effect (see Table 3). The sign of these contributions can be justified, if we realize that the production of the quark sea is associated with a spin flip process (see eq. 30) and hence, the sea quarks has to have an opposite effect. An earlier study highlights the impact of Goldstone Boson (GB) degrees in chiral quark model ($\chi$QM), where the magnetic moments calculated for $p^{*}$ and $n^{*}$ in vacuum, by considering GBE interaction, are 1.4 $\mu_{N}$ and $-0.9\mu_{N}$ for $\chi$QM, which are slightly lower than the corresponding values, 1.7 $\mu_{N}$ and $-1.1\mu_{N}$ as obtained from CQM framework liu . In addition, the magnetic moments procured from $\chi$QM by considering Goldstone Boson Exchange (GBE) and One Gluon Exchange (OGE), separately, gives similar magnetic moment values and are in close agreement to our vacuum state results liu . Two other studies in the literature, with quasibound state approach quasibound and chiral unitary model chiral_unitary have previously calculated for magnetic moments of these two resonances and could be compared from table 7. Table 7 facilitates a comparitive study of the magnetic moments of $p^{*}$ and $n^{*}$ resonances obtained from various models.
Figure 8 depicts the in-medium modifications of $\mu^{*}_{\Sigma^{*+}}$, $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$, as a function of $\rho_{B}/\rho_{0}$ and the behavior of the individual contributions have been plotted in fig. 9. In nuclear medium, the magnetic moments of $\Sigma^{*+}$ initially increases up to $\rho_{B}=2\rho_{0}$ and then saturates at higher densities of the medium. By taking into account the individual contributions from fig. 9, we see that $\mu^{*}_{val}$ increases with density and it has a dominating effect over the cumulative depletion of $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$ contributions. In case of $\Sigma^{*-}$ (at $i_{a}=0,f_{s}=0$), the effective magnetic moment $\mu^{*}_{\Sigma^{*-}}$ decreases and gradually approaches a saturation at very high densities of $\rho_{B}\approx 5\rho_{0}$. Although the in-medium behavior of $\mu^{*}_{val}$ is accompanied by $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$, which imparts opposite contributions, but the valence quarks have a predominant role in determining the nature of $\mu^{*}_{\Sigma^{*-}}$. These opposite contributions due to sea quarks and the orbital motion of sea quarks at higher density regime, evidently explains the stagnant in-medium nature of $\mu^{*}_{\Sigma^{*-}}$ at large medium densities. Further, the density dependent medium modification of $\mu^{*}_{\Sigma^{*0}}$, in symmetric nuclear matter, is much alike the in-medium nature of $\mu^{*}_{\Sigma^{*+}}$ except for a slightly decreasing slope at densities above $1.5\rho_{0}$. The dominance of $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$ contributions over the $\mu^{*}_{sea}$ contribution, justifies the in-medium curve of $\mu^{*}_{\Sigma^{*0}}$ in a nuclear medium. From a point of comparison if you see Table 7, the magnetic moments of $\mu^{*}_{\Sigma^{*+}}$ (1.647 $\mu_{N}$), $\mu^{*}_{\Sigma^{*-}}$ $(-0.754\mu_{N})$ and $\mu^{*}_{\Sigma^{*0}}$ (0.446 $\mu_{N}$) in free space are found to be in agreement with the NCQM results 1.814 $\mu_{N}$, $-0.689\mu_{N}$ and 0.820 $\mu_{N}$, respectively. As seen from Tables 3 and 5, for $f_{s}=0(0.3)$ at $\rho_{B}=3\rho_{0}$, $\mu^{*}_{\Sigma^{*+}}$ and $\mu^{*}_{\Sigma^{*-}}$ values attain 1.697(1.723) $\mu_{N}$ and $-0.770(-0.779)\mu_{N}$. This suggests that as we introduce finite strangeness fraction to the non-strange medium, there will be an enhancement in the nature of in-medium variations of baryonic magnetic moments of $\Sigma^{*+}$, $\Sigma^{*0}$ and $\Sigma^{*-}$. For finite $f_{s}$, the magnetic moments of $\Sigma^{*+}$ and $\Sigma^{*0}$ surges while that of $\Sigma^{*-}$ approaches lower values, instead of attaining a saturation at large densities. For an instance, the in-medium values of $\mu^{*}_{\Sigma^{*+}}$, $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$ (in units of $\mu_{N}$) at $i_{a}=0(0.5)$, when referred to Tables 3 and 4, are $1.697(1.698),-0.770(-0.763)and0.463(0.472)$ in a medium density of $3\rho_{0}$. This and a glance into fig. 8 makes it apparent that the impact of non-zero $i_{a}$, a fine splitting in effective baryonic magnetic moments, is more visible for $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$ as compared to $\mu^{*}_{\Sigma^{*+}}$. Such $i_{a}$ dependent fine shifts in the baryonic magnetic moment have been previously found in works, in the present formalism, dedicated to octet baryons fiftythree and decuplet baryons fiftyfour .
@@
We find it interesting to note that for $\Xi^{*}$ baryons the density effects on $\mu^{*}_{\Xi^{*-}}$ and $\mu^{*}_{\Xi^{*0}}$ values show impressive similarity, owing to the valence quark spin polarizations and corresponding orbital moments of two $s$ quarks that suppresses the impact of light $u$ and $d$ quarks in $\Xi^{*-}$ and $\Xi^{*0}$, respectively sixtytwo . The baryonic magnetic moment of $\Xi^{*-}$, as seen from fig. 10, rises steeply with increase in the relative density ($\rho_{B}/\rho_{0}$) of the medium. Figure 11 clearly illustrates that for $\Xi^{*-}$, the cumulative increasing behavior of $\mu^{*}_{val}$ and $\mu^{*}_{sea}$ contributions overshadows the decreasing orbital contributions from sea quarks ($\mu^{*}_{orbit}$). In a similar manner, a quantitative analysis of fig. 10 and fig. 11 suggests that the density dependent increase in $\mu^{*}_{\Xi^{*0}}$ is well justified by the contributions coming from $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$, which suppresses the contributions coming from $\mu^{*}_{sea}$. Much alike other baryons, we have observed a correlation between finite isospin asymmetry of the medium and a distinguishable splitting in the baryonic magnetic moments in case of $\Xi^{*}$ resonances. Now in order to understand the temperature effects on the in-medium baryonic magnetic moments, we shall lay emphasis on the case of $\mu^{*}_{\Xi^{*-}}$. A close look at fig. 10, shows that for symmetric nuclear matter at medium densities of $\rho_{0}$ and $3\rho_{0}$, at $T=0(100)$ MeV the $\mu^{*}_{\Xi^{*-}}$ values obtained, in $\mu_{N}$ units, are 0.313(0.286) and 0.179(0.163), respectively. This reveals the in-medium nature of $\mu^{*}_{\Xi^{*-}}$, an increasing curve as a function of medium density, is subject to slight reduction in its value as the temperature rises to $T=100$ MeV. In addition, the vacuum state magnetic moments of $\Xi^{*-}$ resonances are found to be much lower than those procured from NCQM (see Table 7).
@@
From fig. 12, we observe that the magnetic moments of ${S^{+}_{11}}(1650)$ and ${S^{0}_{11}}(1650)$ increases, overall as a function of density of the hadronic matter $\rho_{B}$ (in $\rho_{0}$ units). As from fig. 13 it is seen that, the $\mu^{*}_{sea}$ exercises control over the $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$ contributions, which explains the medium modifications of the magnetic moment of ${S^{+}_{11}}(1650)$ in lower density regime (i.e. up till $\rho_{B}=2\rho_{0}$). However, as the nuclear medium grows denser the $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$ contributions demonstrates a dominating influence on the total effective magnetic moment of ${S^{+}_{11}}(1650)$. From fig. 13 it is also observed that for the $N^{*}$ resonance ${S^{0}_{11}}(1650)$, the value of $\mu_{{S^{0}_{11}}(1650)}$ receives major contributions, monotonously, due to the dominant valence quark contribution ($\mu^{*}_{val}$) decreasing with rise in density. For a comparative study based on different models, similar justifications as discussed in the case of $p^{*}$ and $n^{*}$ resonances suffices and the results have been tabulated in Table 7.

### v2

@@ 3.2 In-medium magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances
Figure 6: Variation of effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ (in $\mu_{N}$ units) with baryon density $\rho_{B}$ (in units of $\rho_{0}$).
@@
Figure 7: In-medium trends in explicit contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark ($\mu_{orbit}^{*}$) are presented.
@@
Figure 8: Variation of effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ (in $\mu_{N}$ units) with baryon density $\rho_{B}$ (in units of $\rho_{0}$).
@@
Figure 9: In-medium trends in explicit contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark ($\mu_{orbit}^{*}$) are presented.
@@
Figure 10: Variation of effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ (in $\mu_{N}$ units) with baryon density $\rho_{B}$ (in units of $\rho_{0}$).
@@
Figure 11: In-medium trends in explicit contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark ($\mu_{orbit}^{*}$) are presented.
@@
Figure 12: Variation of effective magnetic moments of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ (in $\mu_{N}$ units) with baryon density $\rho_{B}$ (in units of $\rho_{0}$).
@@
Figure 13: In-medium trends in explicit contributions to effective magnetic moments of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark ($\mu_{orbit}^{*}$) are presented.
@@
This is where we elaborately examine the results of the vacuum and effective magnetic moments of ${\frac{1}{2}}^{-}$ octet baryon resonances $p^{*},n^{*},\Sigma^{*+},\Sigma^{*-},\Sigma^{*0},\Xi^{*-}$, $\Xi^{*0}$ and low-lying $N^{*}$ resonances ${S^{+}_{11}}(1650)$ and ${S^{0}_{11}}(1650)$, observed from figs. 6, 7, 8, 9, 10, 11, 12 and 13. For the sake of discussion, we have also included the individual contributions from valence and sea quarks to the total baryonic magnetic moment which allows us to better interpret the results.
@@
In fig. 6, we have shown the magnetic moment of $p^{*}$ and $n^{*}$ resonances with density of the medium at temperatures $T=0$ MeV and $T=100$ MeV. The observed in-medium behavior of $\mu^{*}_{p^{*}}$ and $\mu^{*}_{n^{*}}$, shows a steeper slope at lower densities of nuclear matter ($\rho_{B}\leq 3\rho_{0}$) and gradually attains saturation at much higher densities. As we trace the plot of these effective baryon magnetic moments (at $T=0$) while the symmetric medium density rises to $\rho_{B}=3\rho_{0}$, we observe that $\mu^{*}_{p^{*}}$ increases by $0.19\mu_{N}$ and $\mu^{*}_{n^{*}}$ initially approaches a minima at $-1.382\mu_{N}$ (at $\rho_{B}\approx 1.1\rho_{0}$) and then climbs up close to its vacuum value at $-1.376\mu_{N}$. A conclusive understanding of the observed effective baryonic magnetic moments of $p^{*}$ and $n^{*}$ can be provided by analyzing fig. 7, where the contributions from valence quarks, sea quarks and orbital angular momentum of the quark sea have been shown individually. For instance, take $n^{*}$ in symmetric nuclear matter, the decrease in the valence quark contributions dominates over the increasing contributions of sea quarks and the orbital contributions caused due to them, which leads to an overall decreasing behavior of $\mu^{*}_{n^{*}}$ below the saturation density of the medium ($\rho_{0}$). As the medium density moves past this point ($\rho_{B}\approx\rho_{0}$), owing to the dominance of $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$ contributions, the effective behavior of $\mu^{*}_{n^{*}}$ reverses and it approaches higher values. On a similar note, the density dependency of $\mu^{*}_{p^{*}}$ can be comprehended by taking into account the increase in $\mu^{*}_{val}$ dominating over the decrease in $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$. In symmetric strange matter at $f_{s}=0.3$, the in-medium behavior of $\mu^{*}_{p^{*}}$ and $\mu^{*}_{n^{*}}$ is enhanced. From fig. 6 we can infer that, the finite strangeness fraction in the medium leads to a splitting in baryonic magnetic moments of $p^{*}$ and $n^{*}$, which is visible beyond the nuclear saturation density, $\rho_{0}$. All other medium parameters affixed, non-zero isospin asymmetry ($i_{a}=0,0.3,0.5$) causes fine splitting in the $\mu^{*}_{p^{*}}$ and $\mu^{*}_{n^{*}}$ curves as can be seen from fig. 6, which has also been identified in their respective $\mu^{*}_{val}$, $\mu^{*}_{sea}$, $\mu^{*}_{orbit}$ contributions by referring to fig. 7. As we have included “Nambu Goldstone Boson Exchange”(NGBE) effect in this study, it is critical to note that the influence of sea quark leads to an opposite sign as compared to the valence quark effect, while on the other hand the orbital angular momentum of sea quarks has the similar contribution as that caused by the valence quarks (see Table 4). The sign of these contributions can be justified, if we realize that the production of the quark sea is associated with a spin flip process (see eq. 32) and hence, the sea quarks has to have an opposite effect. An earlier study highlights the impact of Nambu-Goldstone Boson (NGB) degrees in chiral quark model ($\chi$QM), where the magnetic moments calculated for $p^{*}$ and $n^{*}$ in vacuum, by considering NGBE interaction, are 1.4 $\mu_{N}$ and $-0.9\mu_{N}$ for $\chi$QM, which are slightly lower than the corresponding values, 1.7 $\mu_{N}$ and $-1.1\mu_{N}$ as obtained from CQM framework liu . In addition, the magnetic moments procured from $\chi$QM by considering Nambu Goldstone Boson Exchange (NGBE) and One Gluon Exchange (OGE), separately, gives similar magnetic moment values and are in close agreement to our vacuum state results liu . Two other studies in the literature, with quasibound state approach quasibound and chiral unitary model chiral_unitary have previously calculated for magnetic moments of these two resonances and could be compared from table 8. Table 8 facilitates a comparative analysis of the magnetic moments of $p^{*}$ and $n^{*}$ resonances obtained from various models. Figure 8 depicts the in-medium modifications of $\mu^{*}_{\Sigma^{*+}}$, $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$, as a function of $\rho_{B}/\rho_{0}$ and the behavior of the individual contributions have been plotted in fig. 9. In nuclear medium, the magnetic moments of $\Sigma^{*+}$ initially increases up to $\rho_{B}=2\rho_{0}$ and then saturates at higher densities of the medium. By taking into account the individual contributions from fig. 9, we see that $\mu^{*}_{val}$ increases with density and it has a dominating effect over the cumulative depletion of $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$ contributions. In case of $\Sigma^{*-}$ (at $i_{a}=0,f_{s}=0$), the effective magnetic moment $\mu^{*}_{\Sigma^{*-}}$ decreases and gradually approaches a saturation at very high densities of $\rho_{B}\approx 5\rho_{0}$. Although the in-medium behavior of $\mu^{*}_{val}$ is accompanied by $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$, which imparts opposite contributions, but the valence quarks have a predominant role in determining the nature of $\mu^{*}_{\Sigma^{*-}}$. These opposite contributions due to sea quarks and the orbital motion of sea quarks at higher density regime, evidently explains the stagnant in-medium nature of $\mu^{*}_{\Sigma^{*-}}$ at large medium densities. Further, the density dependent medium modification of $\mu^{*}_{\Sigma^{*0}}$, in symmetric nuclear matter, is much alike the in-medium nature of $\mu^{*}_{\Sigma^{*+}}$ except for a slightly decreasing slope at densities above $1.5\rho_{0}$. The dominance of $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$ contributions over the $\mu^{*}_{sea}$ contribution, justifies the in-medium curve of $\mu^{*}_{\Sigma^{*0}}$ in a nuclear medium. From a point of comparison if you see Table 8, the magnetic moments of $\mu^{*}_{\Sigma^{*+}}$ (1.647 $\mu_{N}$), $\mu^{*}_{\Sigma^{*-}}$ $(-0.754\mu_{N})$ and $\mu^{*}_{\Sigma^{*0}}$ (0.446 $\mu_{N}$) in free space are found to be in agreement with the NCQM results 1.814 $\mu_{N}$, $-0.689\mu_{N}$ and 0.820 $\mu_{N}$, respectively. As seen from Tables 4 and 6, for $f_{s}=0(0.3)$ at $\rho_{B}=3\rho_{0}$, $\mu^{*}_{\Sigma^{*+}}$ and $\mu^{*}_{\Sigma^{*-}}$ values attain 1.697(1.723) $\mu_{N}$ and $-0.770(-0.779)\mu_{N}$. This suggests that as we introduce finite strangeness fraction to the non-strange medium, there will be an enhancement in the nature of in-medium variations of baryonic magnetic moments of $\Sigma^{*+}$, $\Sigma^{*0}$ and $\Sigma^{*-}$. For finite $f_{s}$, the magnetic moments of $\Sigma^{*+}$ and $\Sigma^{*0}$ surges while that of $\Sigma^{*-}$ approaches lower values, instead of attaining a saturation at large densities. For an instance, the in-medium values of $\mu^{*}_{\Sigma^{*+}}$, $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$ (in units of $\mu_{N}$) at $i_{a}=0(0.5)$, when referred to Tables 4 and 5, are $1.697(1.698),-0.770(-0.763)and0.463(0.472)$ in a medium density of $3\rho_{0}$. This and a glance into fig. 8 makes it apparent that the impact of non-zero $i_{a}$, a fine splitting in effective baryonic magnetic moments, is more visible for $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$ as compared to $\mu^{*}_{\Sigma^{*+}}$. Such $i_{a}$ dependent fine shifts in the baryonic magnetic moment have been previously found in works, in the present formalism, dedicated to octet baryons fiftythree and decuplet baryons fiftyfour .
We find it interesting to note that for $\Xi^{*}$ baryons the density effects on $\mu^{*}_{\Xi^{*-}}$ and $\mu^{*}_{\Xi^{*0}}$ values show impressive similarity, owing to the valence quark spin polarizations and corresponding orbital moments of two $s$ quarks that suppresses the impact of light quarks present in them sixtytwo .
@@
The baryonic magnetic moment of $\Xi^{*-}$, as seen from fig. 10, rises steeply with increasing relative hadronic matter density ($\rho_{B}/\rho_{0}$). Figure 11 clearly illustrates that for $\Xi^{*-}$, the cumulative increasing behavior of $\mu^{*}_{val}$ and $\mu^{*}_{sea}$ contributions overshadows the decreasing orbital contributions from sea quarks ($\mu^{*}_{orbit}$). In a similar manner, a quantitative analysis of fig. 10 and fig. 11 suggests that the density dependent increase in $\mu^{*}_{\Xi^{*0}}$ is well justified by the contributions coming from $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$, which suppresses the contributions coming from $\mu^{*}_{sea}$. Much alike other baryons, we have observed a correlation between finite isospin asymmetry of the medium and a distinguishable splitting in the baryonic magnetic moments in case of $\Xi^{*}$ resonances. Now in order to understand the temperature effects on the in-medium baryonic magnetic moments, we shall lay emphasis on the case of $\mu^{*}_{\Xi^{*-}}$. A close look at fig. 10, shows that for symmetric nuclear matter at medium densities of $\rho_{0}$ and $3\rho_{0}$, at $T=0(100)$ MeV the $\mu^{*}_{\Xi^{*-}}$ values obtained, in $\mu_{N}$ units, are 0.313(0.286) and 0.179(0.163), respectively. This reveals the in-medium nature of $\mu^{*}_{\Xi^{*-}}$, an increasing curve as a function of medium density, is subject to slight reduction in its value as the temperature rises to $T=100$ MeV. In addition, the vacuum state magnetic moments of $\Xi^{*-}$ resonances are found to be much lower than those procured from NCQM (see Table 8).
@@
From fig. 12, we observe that the magnetic moments of ${S^{+}_{11}}(1650)$ and ${S^{0}_{11}}(1650)$ increases, overall as a function of relative medium density $\rho_{B}/\rho_{0}$. As from fig. 13 it is seen that, the $\mu^{*}_{sea}$ exercises control over the $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$ contributions, which explains the medium modifications of the magnetic moment of ${S^{+}_{11}}(1650)$ in lower density regime (i.e. up till $\rho_{B}=2\rho_{0}$). However, as the nuclear medium grows denser the $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$ contributions demonstrates a dominating influence on the total effective magnetic moment of ${S^{+}_{11}}(1650)$. From fig. 13 it is also observed that for the $N^{*}$ resonance ${S^{0}_{11}}(1650)$, the value of $\mu_{{S^{0}_{11}}(1650)}$ receives major contributions, monotonously, due to the dominant valence quark contribution ($\mu^{*}_{val}$) decreasing with rise in density. For a comparative study based on different models, similar justifications as discussed in the case of $p^{*}$ and $n^{*}$ resonances suffices and the results have been tabulated in Table 8.